OpenStudy (friendzone):

write the following in factored form by factoring out the greatest common factor 66y to the 12th power + 42y to the 8th power=

OpenStudy (anonymous):

\[66y^{12}+42y^8\] right?

OpenStudy (friendzone):

yes

OpenStudy (anonymous):

the greatest common divisor of 66 and 42 is 6

OpenStudy (anonymous):

ie. you can factor a 6 out of both terms also you can factor out a \(y^8\)

OpenStudy (friendzone):

correct you can get alot out of it

OpenStudy (anonymous):

\[66y^{12}+42y^8\] factor the 6 and \(y^8\) get \[6y^8(11y^4+7)\]

OpenStudy (igreen):

Yep, since the greatest common factor is \(6y^8\), you can divide that to both terms. \(66y^{12} \div 6y^8 = 11^4\) - Divide the whole numbers, subtract the exponents. \(42y^8 \div 6y^8 = 7\) - Same thing here.

OpenStudy (igreen):

Then we just put them in a parenthesis, with the greatest common factor outside of it. \(6y^8(11y^4 + 7)\)

OpenStudy (igreen):

*\(11y^4\)

OpenStudy (friendzone):

good help thanks so much